# BitMark for Infinity: Watermarking Bitwise Autoregressive Image Generative Models

### v1

@@ Abstract
State-of-the-art text-to-image models like Infinity generate photorealistic images at an unprecedented speed. These models operate in a bitwise autoregressive manner over a discrete set of tokens that is practically infinite in size. However, their impressive generative power comes with a growing risk: as their outputs increasingly populate the Internet, they are likely to be scraped and reused as training data—potentially by the very same models. This phenomenon has been shown to lead to model collapse, where repeated training on generated content, especially from the models’ own previous versions, causes a gradual degradation in performance. A promising mitigation strategy is watermarking, which embeds human-imperceptible yet detectable signals into generated images—enabling the identification of generated content. In this work, we introduce BitMark, a robust bitwise watermarking framework for Infinity. Our method embeds a watermark directly at the bit level of the token stream across multiple scales (also referred to as resolutions) during Infinity’s image generation process. Our bitwise watermark subtly influences the bits to preserve visual fidelity and generation speed while remaining robust against a spectrum of removal techniques. Furthermore, it exhibits high radioactivity, i.e., when watermarked generated images are used to train another image generative model, this second model’s outputs will also carry the watermark. The radioactive traces remain detectable even when only fine-tuning diffusion or image autoregressive models on images watermarked with our BitMark. Overall, our approach provides a principled step toward preventing model collapse in image generative models by enabling reliable detection of generated outputs.
@@ 1 Introduction
The image generation capabilities of large scale models [7, 10, 21, 24] have improved tremendously. These models can now generate at an unprecedented speed (of less than a second) high-quality photorealistic outputs, which are often indistinguishable from real images [10]. One of the current state-of-the-art image generative models is Infinity [10], which operates in a bitwise autoregressive manner [32, 37] over an almost infinite number of tokens and yields high-resolution images. This performance, however, comes at the expense of a training process that is highly time-, cost- and most of all data-intensive [23, 27]. The required training data are often scraped from the Internet [11, 22, 30]. This practice carries risks: with the ever-growing amount of generated data on the internet, the models get increasingly trained on generated data rather than natural data. The iterative training of models on their own generated data has been shown to degrade performance [1, 31], which is referred to as model collapse. The process also leads to encoding the current models’ mistakes, biases, and unfairnesses into its future versions [36]. From the perspective of model owners, especially big companies such as OpenAI, Google, or ByteDance, who invest in large scale proprietary models that are exposed via public APIs, this poses a significant threat. To prevent the possible degradation of model quality, it is crucial for them to identify their own generated content.
@@
To close this gap and provide a robust way of detecting images generated by state-of-the-art models like Infinity [10], we propose the BitMark watermark that aligns with Infinity’s bitwise image generation process and operates directly on a bit-level. We show that this bit-level design is crucial in image autoregressive models, because watermarks at the token level—a common practice in autoregressive language models [14, 41]—are suboptimal. This is because unlike language models, which operate on discrete tokens and yield stable token representations, image models must decode to and re-encode from a continuous image space, introducing discrepancies between the original and recovered tokens. These discrepancies can decrease the watermark signal. Yet, we find that the discrepancies are substantially smaller at the bit level (Figure 1), motivating our bitwise approach to obtain a robust watermark. To embed the watermark, BitMark first divides all possible n-bit sequences into green and red lists, where the green list is supposed to contain sequences indicating the presence of the watermark. The embedding process of the watermark aims at completing one of the green-list bit sequences for every next bit by slightly increasing each logit of the preferred bit that would complete the sequence. This eventually causes the generated image to contain more sequences from the green list than natural images, which can be used for detection. Concretely, the detection of the watermark is performed on re-encoded images and based on the count of bit sequences from the green list.
Our watermark exhibits many desirable properties. Most importantly, the generated images are of the same quality as images generated without the embedded watermark. Moreover, the watermark embedding has a negligible impact on the speed of the image generation. Also, the corresponding watermark detection method is fast and reliable. As in natural images all bit sequences are equally likely to occur while our watermarked images have the bias towards sequences from the green list, natural images are extremely unlikely to be detected as watermarked. Furthermore, working on the bit-level makes BitMark robust against a wide range of attacks, such as adding Gaussian noise or blur to the image, applying color jitter, random crop, or JPEG compression. Our watermark is also highly resilient against dedicated watermark-removal techniques, for example, the watermark in the sand attack [38] or CtrlRegen [18]. Finally, we show that BitMark is radioactive. This means that when we use the watermarked images to train or fine-tune other image generative models, our watermark is still present in the outputs from these other models. We demonstrate all these properties through our extensive experimental evaluation.
@@
  • We propose BitMark—the first watermarking scheme for Infinity. BitMark leverages Infinity’s unique structure and bitwise image generation process to embed the watermark on the level of bits, which preserves the high quality of the models’ outputs and its inference speed. Our approach can be applied at inference time and does not require additional retraining.
@@ 2 Background
We consider the image autoregressive models, with the focus on the state-of-the-art text-to-image Infinity model that operates in a autoregressive manner, and the techniques used to robustly watermark image generative models.
Visual AutoRegressive Modeling (VAR) [32]. VAR leverages discrete quantization to predict tokens as residual maps, scaling from an initial low scale (coarse-grain) to a final scale (fine-grain) resolution. VAR further leverages the transformer architecture [33] across all tokens to significantly improve generation quality. VAR was further extended from class-to-image to the Infinity model [10], which is a text-to-image autoregressive model for high-resolution image synthesis, using next-scale prediction similar to VAR [32]. Instead of leveraging index-wise tokens for quantization like VAR [32], Infinity uses bitwise token prediction with an implicit codebook and applies binary spherical quantization [42] to select an entry from the codebook. This results in more efficient computation and further enables the usage of large codebook sizes, e.g., $2^{64}$. Random bit flipping further empowers the model to self-correct on each scale, leveling out errors stemming from earlier tokens. The visual autoregressive paradigm instantiated by Infinity achieves the state-of-the-art (SOTA) in image generation.
@@
Watermarking Autoregressive Models. As of now, all the efforts on watermarking autoregressive models have been directed towards creating watermarks for Large Language Models (LLMs) [14]. The KGW watermark [14] uses the concept of a green and red token lists, where generation is softly nudged to predict tokens from the green list rather than from the red list. The lists are generated dynamically, based on the hash value of the previously predicted token(s). For detection, the LLM used to generate the text is not needed. Instead, a suspect text is assessed based on how often tokens from the green list appear, compared to tokens from the red list. At the end of detection a statistical test is performed. KGW was extended to the Unigram-Watermark [41], which simplifies the design by introducing a fixed global green and red lists. The simplified approached allowed for guaranteed generation quality, provable correctness in watermark detection, and robustness against text editing and paraphrasing. In this work, we we build on these watermarking schemes originally developed for large language models (LLMs) to address the lack of watermarking techniques for image autoregressive models, with focus on the state-of-the-art Infinity model.
@@ 3 BitMark Method
The Infinity architecture [10] follows an autoregressive generation paradigm, but instead of directly predicting tokens like other autoregressive models, such as VAR [32] or RAR [37], it predicts bits. We leverage this prediction scheme to design a dedicated watermark. Our BitMark watermark is embedded across all scales (resolutions) of Infinity, which subtly shift the bitwise representation towards a higher frequency of detectable patterns to resist watermark removal and manipulation.
Problem Formulation and Desiderata. Our goal is to embed a watermark during Infinity’s image generation process and subsequently detect it by analyzing the generated image. BitMark targets the prevention of model collapse. In this setup, only the model owner is concerned with the watermark and wants to ensure that their own data are not used to train their next models. Thus, we aim at four essential properties for our watermarking scheme to prevent model collapse: (1) Negligible Impact on the Generation. To maintain the utility of the generative model, the watermark should not degrade the quality of the generated images, while also maintaining a fast inference speed. (2) Reliable Verification. The model owner should be able to detect the watermarked images and verify the confidence of the extracted watermark. The model owner encodes the image to their own space and it is only relevant to the model owner if the watermark is present there or not. The encoder part of the model is also kept private so only the model owner can verify the watermark. The natural images should not be detected as containing the watermark. At the scale of the large models trained on trillions of samples, a small number of false positives is acceptable as long as it does not exceeds a certain threshold (e.g., 1%), which can also prevent model poisoning [40]. (3) Robustness. The watermark should be robust not only against conventional attacks, such as Gaussian noise or JPEG compression, but also against strong dedicated watermark removal attacks such as CtrlRegen [18] or watermarks in the sand [38]. (4) Radioactivity. If other generative models are trained on watermarked images, images generated by these new models should also contain our watermarks [6, 26, 29].
@@ 3.1 Motivation for the Bitwise Watermark
We identify an inherent challenge in watermarking the images generated by autoregressive models like Infinity, stemming from imperfections in image encoding and decoding. Specifically, the model first generates a sequence of tokens, which are decoded into an image. However, when the exact same image is re-encoded and tokenized, it produces a different set of tokens.
@@
This also holds partly true for bits that underlie the tokens. The observed behavior does not occur in the autoregressive LLMs, where the tokens produced by decoding and then re-encoding a generated text match exactly. We quantify the bit and token discrepancies in Figure 1. In both cases, bits and tokens are generated by Infinity and then reconstructed by its auto-encoder. For Infinity, the highest token overlap occurs at the initial and final resolutions (scales). This is likely because the early stages capture coarse structural features, while the final resolution refines the image, leading to fewer changes in the token sequence. We observe a consistent average token overlap of approximately 2.385% and the bit overlap of 77.432%. Given the significantly large codebook size of $2^{n}$ (where $n\in\{32,48,64\}$) in Infinity, we observe that the bit-wise information loss is much smaller than the token-wise loss. This motivates our work to design a watermark that biases the generated sequences of tokens on the bit-level instead of on the token-level.
@@ 3.2 Embedding our BitMark Watermark
Next, we detail how BitMark is embedded. The embedding protocol is directly applied on the bit-level in the autoregressive generation process of the Infinity model. We present the full procedure in Algorithm 1. First, we separate the list of all bit sequences of length $n$: $S=\{0,1\}^{n}$ into a red list $R$ and a green list $G$. Intuitively, we softly nudge the model to generate bit sequences of length $n$ that are part of the green list $G$. For each next bit prediction $b_{j}$, if the bit can complete one of the bit sequences from $G$, we add a small bias $\delta$ to the logit $l_{j}^{(b_{j})}$ of the bit $b_{j}$:
@@
where $pre=(b_{j-(n-1)},\dots,b_{j-1})$ is the prefix bit sequence of length $(n-1)$, i.e., the sequence of bits preceding $b_{j}$, $+$ is the bit concatenation operation, and $l_{j}$ is the vector of logits for the bit $b_{j}$. We apply this procedure to all bits and across all resolutions (scales) that are generated by the model. Thus, as many bit sequences as possible are biased towards $G$ and the final green fraction, i.e., the number of sequences part of the green set, of the generated bits is significantly above the threshold of 50%, the standard fraction for clean images (see Appendix D). By biasing the logit for the preferred bit, BitMark has negligible impact on image quality, as mostly high entropy bits are flipped (see Table 1). Since high entropy bits have sufficiently small difference in the probability of assigning them to value 00 or $1$, the relatively small added bias influences the final value. Additionally, biasing only the logits makes the watermarking process difficult to revert, as it is difficult to know for an attacker which bits are of high entropy and can be flipped to the red list $R$ without degrading the final image quality.
@@ 3.3 Watermark Detection
Detecting the Watermark. The detection protocol follows the standard encoding from Infinity and is detailed in Algorithm 2. First we obtain the bit representations for each resolution. We then use our knowledge of $G$ and $R$ to test how often the bit sequences from the red and green lists appear and count them. This results in the knowledge of how many green sequences appear in the encoded image, which is random for every non-watermarked image, i.e., $50\%$ (see Appendix D).
Robust Statistical Testing. Given the final number of green and red sequences detected in the encoded image, we perform a statistical test based on the null hypothesis $\mathcal{H}_{0}:$ The sequence of bits is generated with no knowledge about the green and red lists. Intuitively, this assumes that a given image is not watermarked. A watermarked image consists of $\gamma T$ green and $(1-\gamma)T$ red bit sequences, where $\gamma$ determines the size of the green list and $T$ is the total number of bits in the encoded image that our watermark can influence. Per a given token $t$ with $m$ bits, BitMark influences $k=m-(n-1)$ bits, since we start watermarking from the first $n$ bit pattern. Thus, overall we have $T=\sum_{i=1}^{K}r_{i}\cdot k$, where $r_{i}$ is number of tokens per resolution $i$. A natural image is expected to violate the red list rule with the half of its bits, while a watermarked image violates this rule less frequently. If the null hypothesis is true, then the number of green list bit sequences denoted C has the expected value of $\gamma T$ and variance $\gamma(1-\gamma)T$. To ensure statistical validity, we follow [14] and use a z-test. We reject the null hypothesis and detect the watermark if $z$ is above a chosen threshold:
$z=(C-\gamma T)/\sqrt{T\gamma(1-\gamma)}.$ (2)
@@
Inputs: prompt $x$, tokenizer $tok$, autoregressive model $f_{\theta}$ with parameters $\theta$, green set $G$, red list $R$, image decoder $\mathcal{D}$, BSQ quantizer $\mathcal{Q}$.
Hyperparameters: steps $K$ (number of resolutions), resolutions $(h_{i},w_{i})_{i=1}^{K}$, the number of tokens for resolution $i$ is $r_{i}$, constant $\delta$ added to the logits of the $green$ list bits, $n$ the length of the bit sequences in the lists ($G$ and $R$), the number of bits per token $d$.
$e=tok(x)$
@@
       $(l_{1},\dots,l_{r_{i}\cdot d})=f_{\theta}(e,h_{i},w_{i})$for $t=1,\dots,r_{i}$ do
             for $j=n,\dots,d$ do
[…]
      $u_{i}=(b_{1},\dots,b_{r_{i}\cdot d})$$z_{i}=\mathcal{Q}(u_{i})$ $z_{i}=\text{upsample}(z_{i},h_{K},w_{K})$$e=e+z_{i}$
@@
Return: watermarked image $im$
@@
Inputs: raw image $im$, green list $G$, red list $R$, image encoder $\mathcal{E}$, BSQ quantizer $\mathcal{Q}$
Hyperparameters: steps $K$ (number of resolutions), resolutions $(h_{i},w_{i})_{i=1}^{K}$, the number of tokens for resolution $i$ is $r_{i}$, the number of bits per token $d$.
@@
       $u_{i}=\mathcal{Q}(\text{downsample}(e,h_{i},w_{i}))$ $u_{i}=(b_{1},\dots,b_{r_{i}\cdot d})$ $C$ = Count($(b_{1},\dots,b_{r_{i}\cdot d})$, $G$) $u_{i}=\text{upsample}(u_{i},h_{K},w_{K})$$e=e-u_{i}$
Return: StatisticalTest($\mathcal{H}_{0},(C)$)

### v2

# BitMark: Watermarking Bitwise Autoregressive Image Generative Models

@@ Abstract
State-of-the-art text-to-image models generate photorealistic images at an unprecedented speed. This work focuses on models that operate in a bitwise autoregressive manner over a discrete set of tokens that is practically infinite in size. However, their impressive generative power comes with a growing risk: as their outputs increasingly populate the Internet, they are likely to be scraped and reused as training data—potentially by the very same models. This phenomenon has been shown to lead to model collapse, where repeated training on generated content, especially from the models’ own previous versions, causes a gradual degradation in performance. A promising mitigation strategy is watermarking, which embeds human-imperceptible yet detectable signals into generated images—enabling the identification of generated content. In this work, we introduce BitMark, a robust bitwise watermarking framework. Our method embeds a watermark directly at the bit level of the token stream during the image generation process. Our bitwise watermark subtly influences the bits to preserve visual fidelity and generation speed while remaining robust against a spectrum of removal techniques. Furthermore, it exhibits high radioactivity, i.e., when watermarked generated images are used to train another image generative model, this second model’s outputs will also carry the watermark. The radioactive traces remain detectable even when only fine-tuning diffusion or image autoregressive models on images watermarked with our BitMark. Overall, our approach provides a principled step toward preventing model collapse in image generative models by enabling reliable detection of generated outputs. The code is available at https://github.com/sprintml/BitMark.
@@ 1 Introduction
The image generation capabilities of large scale models [8, 11, 37, 12, 23, 26] have improved tremendously. These models can now generate at an unprecedented speed high-quality photorealistic outputs, which are often indistinguishable from real images [11, 37]. Several models [11, 37] now achieve state-of-the-art performance by operating in a bitwise autoregressive manner [35, 41]. This provides them with a near infinite number of possible tokens, yielding high-resolution images. This performance, however, comes at the expense of a training process that is highly time-, cost- and most of all data-intensive [25, 29]. The required training data are often scraped from the Internet [13, 24, 32]. This practice carries risks: with the ever-growing amount of generated data on the internet, the models get increasingly trained on generated data rather than natural data. The iterative training of models on their own generated data has been shown to degrade performance [1, 33], which is referred to as model collapse. The process also leads to encoding the current models’ mistakes, biases, and unfairnesses into its future versions [40]. From the perspective of model owners, especially big companies such as OpenAI, Google, or ByteDance, who invest in large scale proprietary models that are exposed via public APIs, this poses a significant threat. To prevent the possible degradation of model quality, it is crucial for them to identify their own generated content.
@@
To close this gap and provide a robust way of detecting images generated by state-of-the-art autoregressive models operating on bits like Infinity [11] and Instella [37]we propose the BitMark watermark that aligns with bitwise image generation process and operates directly on a bit-level. We show that this bit-level design is crucial in image autoregressive models, as watermarks at the token level—a common practice in autoregressive language models [16, 46]—are suboptimal. This is because unlike language models, which operate on discrete tokens and yield stable token representations, image models must decode to and re-encode from a continuous image space, introducing discrepancies between the original and recovered tokens. These discrepancies can decrease the watermark signal. Yet, we find that the discrepancies are substantially smaller at the bit level (Figure˜1), motivating our bitwise approach to obtain a robust watermark. To embed the watermark, BitMark first divides all possible n-bit sequences into green and red lists, where the green list is supposed to contain sequences indicating the presence of the watermark. The embedding process of the watermark aims at completing one of the green-list bit sequences for every next bit by slightly increasing each logit of the preferred bit that would complete the sequence. This eventually causes the generated image to contain more sequences from the green list than natural images, which can be used for detection. Concretely, the detection of the watermark is performed on re-encoded images and based on the count of bit sequences from the green list.
Our watermark exhibits many desirable properties. Most importantly, the generated images are of similar quality as images generated without the embedded watermark. Moreover, the watermark embedding has a negligible impact on the speed of the image generation. Also, the corresponding watermark detection method is fast and reliable. As in natural images all bit sequences are equally likely to occur while our watermarked images have the bias towards sequences from the green list, natural images are extremely unlikely to be detected as watermarked. Furthermore, working on the bit-level makes BitMark robust against a wide range of attacks, such as adding Gaussian noise or blur to the image, applying color jitter, random crop, or JPEG compression. Our watermark is also highly resilient against dedicated watermark-removal techniques, for example, the watermark in the sand attack [42] or CtrlRegen [20]. Finally, we show that BitMark is radioactive. This means that when we use the watermarked images to train or fine-tune other image generative models, our watermark is still present in the outputs from these other models. We demonstrate all these properties through our extensive experimental evaluation.
@@
  • We propose BitMark—the first watermarking scheme for bitwise image generative models. BitMark leverages the bitwise image generation process of these models to embed the watermark on the level of bits, which preserves the high quality of the models’ outputs and its inference speed. We show the success of our method on Infinity and Instella.
@@ 2 Background
We first introduce image autoregressive models, with a focus on the state-of-the-art text-to-image models that operate on bits, namely Infinity and Instellaoperating in an autoregressive manner. We then provide an overview on techniques used to robustly watermark image generative models.
Visual AutoRegressive Modeling (VAR) [35]. VAR leverages discrete quantization to predict tokens as residual maps, scaling from an initial low scale (coarse-grain) to a final scale (fine-grain) resolution. VAR further leverages the transformer architecture [36] across all tokens to significantly improve generation quality. VAR was extended from class-to-image to the Infinity model [11], which is a text-to-image autoregressive model for high-resolution image synthesis, using next-scale prediction similar to VAR [35]. Instead of leveraging index-wise tokens for quantization like VAR [35], Infinity uses bitwise token prediction with an implicit codebook and applies binary spherical quantization [47] to select an entry from the codebook. This results in more efficient computation and further enables the usage of large codebook sizes, e.g., $2^{64}$. Random bit flipping further empowers the model to self-correct on each scale, leveling out errors stemming from earlier tokens. The visual autoregressive paradigm instantiated by Infinity achieves the state-of-the-art (SOTA) in image generation. Instella [37] is a token-lightweight next-patch prediction autoregressive model. It efficiently leverages bits to minimize the number of tokens required for generating high-resolution images, reducing it to as few as 128 tokens. This bitwise paradigm allows Instella to achieve competitive results in image generation with significantly lower computational resources.
@@
Watermarking Autoregressive Models. As of now, all the efforts on watermarking autoregressive models have been directed towards creating watermarks for Large Language Models (LLMs) [16]. The KGW watermark [16] uses the concept of a green and red token lists, where generation is softly nudged to predict tokens from the green list rather than from the red list. The lists are generated dynamically, based on the hash value of the previously predicted token(s). For detection, the LLM used to generate the text is not needed. Instead, a suspect text is assessed based on how often tokens from the green list appear, compared to tokens from the red list. At the end of detection a statistical test is performed. KGW was extended to the Unigram-Watermark [46], which simplifies the design by introducing a fixed global green and red lists. The simplified approached allowed for guaranteed generation quality, provable correctness in watermark detection, and robustness against both text editing and paraphrasing. In this work, we build on these watermarking schemes originally developed for large language models to address the lack of watermarking techniques for bitwise generative models.
@@ 3 BitMark Method
The autoregressive image generative models that we consider, namely Infinity [11] and Instella [37], predict bits instead of tokens, distinguishing them from other image autoregressive models such as VAR [35] or RAR [41]. We leverage this prediction scheme to design a dedicated watermark. Our BitMark watermark is embedded during the generation process, subtly shifting the bitwise representation towards a higher frequency of detectable patterns to resist manipulation.
Problem Formulation and Desiderata. Our goal is to embed a watermark during the image generation process and subsequently detect it by analyzing the generated image. BitMark targets the prevention of model collapse. In this setup, only the model owner is concerned of detecting the watermark and wants to ensure that their own data are not used to train their next models. Thus, we aim at four essential properties for our watermarking scheme to prevent model collapse: (1) Negligible Impact on the Generation. To maintain the utility of the generative model, the watermark should not degrade the quality of the generated images, while also maintaining a fast inference speed. (2) Reliable Verification. The model owner should be able to detect the watermarked images and verify the confidence of the extracted watermark. The model owner encodes the image to their own space and it is only relevant to the model owner if the watermark is present there or not. The encoder part of the model is also kept private so only the model owner can verify the watermark. The natural images should not be detected as containing the watermark. At the scale of the large models trained on trillions of samples, a small number of false positives is acceptable as long as it does not exceeds a certain threshold (e.g., 1%), which can also prevent model poisoning [45]. (3) Robustness. The watermark should be robust not only against conventional attacks, such as Gaussian noise or JPEG compression, but also against strong dedicated watermark removal attacks such as CtrlRegen [20] or watermarks in the sand [42]. (4) Radioactivity. If other generative models are trained on watermarked images, images generated by these new models should also contain our watermarks [7, 28, 31].
@@ 3.1 Motivation for the Bitwise Watermark
We identify an inherent challenge in watermarking the images generated by autoregressive models stemming from imperfections in image encoding and decoding. Specifically, the model first generates a sequence of tokens, which are decoded into an image. However, when the exact same image is re-encoded and tokenized, it produces a different set of tokens.
@@
This also holds partly true for bits that underlie the tokens. The observed behavior does not occur in the autoregressive LLMs, where the tokens produced by decoding and then re-encoding a generated text match exactly. We quantify the bit and token discrepancies in Figure˜1 for Infinity. In both cases, bits and tokens are generated by the model and then reconstructed by its auto-encoder. Here, the highest bit overlap occurs at the initial and final resolutions (scales). This is likely because the early stages capture coarse structural features, while the final resolution refines the image, leading to fewer changes in the bit sequence. We observe a consistent average token overlap of approximately 2.385% and bit overlap of 77.432%. Given the significantly large codebook size of $2^{32}$ in Infinity, we observe that the bitwise information loss is much smaller than the token-wise loss. This motivates our work to design a watermark that biases the generated sequences of tokens on the bit-level instead of on the token-level.
@@ 3.2 Embedding our BitMark Watermark
Next, we detail how BitMark is embedded. The embedding protocol is directly applied on the bit-level in the autoregressive generation process of the model. We present the full procedure for multi-scale models like Infinity in Algorithm˜1 and for single-scale models like Instella IAR in Algorithm˜3. First, we separate the list of all bit sequences of length $n$: $S=\{0,1\}^{n}$ into a red list $R$ and a green list $G$. Intuitively, we softly nudge the model to generate bit sequences of length $n$ that are part of the green list $G$. For each next bit prediction $b_{j}$, if the bit can complete one of the bit sequences from $G$, we add a small bias $\delta$ to the logit $l_{j}^{(b_{j})}$ of the bit $b_{j}$:
@@
where $pre=(b_{j-(n-1)},\dots,b_{j-1})$ is the prefix bit sequence of length (n1)(n-1), i.e., the sequence of bits preceding $b_{j}$, ++ is the bit concatenation operation, and $l_{j}$ is the vector of logits for the bit $b_{j}$. We apply this procedure to all bits that are generated by the model. Thus, as many bit sequences as possible are biased towards $G$ and the final green fraction, i.e., the number of sequences part of the green set, of the generated bits is significantly above the threshold of 50%, the standard fraction for clean images (see Section˜F.6). By biasing the logit for the preferred bit, BitMark has negligible impact on image quality, as mostly high entropy bits are flipped (see Table˜1). Since high entropy bits have sufficiently small difference in the probability of assigning them to value 0 or $1$, the relatively small added bias influences the final value. Additionally, biasing only the logits makes the watermarking process difficult to revert, as it is difficult to know for an attacker which bits are of high entropy and can be flipped to the red list $R$ without degrading the final image quality (see Figure˜9).
@@ 3.3 Watermark Detection
Detecting the Watermark. The detection protocol follows the standard encoding of the model and is detailed for multi-scale models in Algorithm˜2 and for single-scale models in Algorithm˜4. First we obtain the bit representations for each resolution. We then use our knowledge of $G$ and $R$ to test how often the bit sequences from the red and green lists appear and count them. This results in the knowledge of how many green sequences appear in the encoded image, which is random for every non-watermarked image, i.e., $50\%$ (see Section˜F.6).
Robust Statistical Testing. Given the final number of green and red sequences detected in the encoded image, we perform a statistical test based on the null hypothesis $\mathcal{H}_{0}:$ The sequence of bits is generated with no knowledge about the green and red lists. Intuitively, this assumes that a given image is not watermarked. A watermarked image consists of $\gamma T$ green and $(1-\gamma)T$ red bit sequences, where $\gamma$ determines the size of the green list and $T$ is the total number of bits in the encoded image that our watermark can influence. Per a given token $t$ with $m$ bits, BitMark influences k=m(n1)k=m-(n-1) bits, since we start watermarking from the first $n$ bit pattern. Thus, overall we have $T=\sum_{i=1}^{K}r_{i}\cdot k$, where $r_{i}$ is number of tokens per resolution $i$. A natural image is expected to violate the red list rule with the half of its bits, while a watermarked image violates this rule less frequently. If the null hypothesis is true, then the number of green list bit sequences denoted C has the expected value of $\gamma T$ and variance $\gamma(1-\gamma)T$. To ensure statistical validity, we follow [16] and use a z-test:
$z=(C-\gamma T)/\sqrt{T\gamma(1-\gamma)}$ (2)
We reject the null hypothesis and detect the watermark if $z$ is above a chosen threshold.
@@
Inputs: autoregressive model $f_{\theta}$ with parameters $\theta$, green list $G$, red list $R$, image decoder $\mathcal{D}$.
Hyperparameters: steps $K$ (number of resolutions), resolutions $(h_{i},w_{i})_{i=1}^{K}$, the number of tokens for resolution $i$ is $r_{i}$, constant $\delta$ added to the logits of the greengreen list bits, $n$ - the length of the bit sequences in the lists ($G$ and $R$), $m$ the number of bits per token.
@@
  $(l_{1},\dots,l_{r_{i}\cdot m})=f_{\theta}(e,h_{i},w_{i})$for $t=1,\dots,r_{i}$ do
     for $j=n,\dots,m$ do
        $c=(t-1)\cdot m+j\text{ //counter}$$p_{c}=\text{Bias}(G,l_{c},\delta)$$s_{c}=\text{Softmax}(p_{c})$$b_{c}=\text{Sample}(s_{c})$
 $u_{i}=(b_{1},\dots,b_{r_{i}\cdot m})$$z_{i}=\text{Lookup}(u_{i})$$z_{i}=\text{Interpolate}(z_{i},h_{K},w_{K})$$e=e+\phi_{i}(z_{i})$
@@
Return: watermarked image imim
@@
Inputs: raw image $im$, green list $G$, red list $R$, image encoder $\mathcal{E}$, quantizer $\mathcal{Q}$
Hyperparameters: steps $K$ (number of resolutions), resolutions $(h_{i},w_{i})_{i=1}^{K}$, the number of tokens for resolution $i$ is $r_{i}$, $n$ - the length of the bit vector.
@@
[…]